OpenStudy (anonymous):

A 10-ft. board is cut into 2 pieces so that one piece is 2 feet longer than 3 times the shorter piece. If the shorter piece is x feet long,find the lengths of both pieces. with steps please

OpenStudy (anonymous):

x+(3x+2)=10 x is the length of the shorter piece.

OpenStudy (anonymous):

Not sure how to solve that..

OpenStudy (anonymous):

x+3x=4x, so 4x+2=10.. Does that help?

OpenStudy (anonymous):

Not really :/

OpenStudy (anonymous):

Subtract the 2, so 4x=8, divide by 4. so x=2.That means the shorter piece is 2 feet, and the longer piece is 8 feet.

OpenStudy (anonymous):

oh ok i got it now thank you so much :)

OpenStudy (anonymous):

No problem:]